An ion X2+(the charge is 2+) contains 18 electron and 20 neutrons.Calculate the atomic number and mass number of X.
1
Expert's answer
2013-03-11T10:13:13-0400
number of neutrons = 20

a dipostive cation has 2 electrons less thus originally the no of electrons = 18 + 2 = 20 = atomic number

no. of protons=no of electrons in an neutral atom,

thus no. of protons = 20

Atomic mass = no. of protons + no. of neutrons

atomic mass = 20 + 20 = 40